# Star Attention: Efficient LLM Inference over Long Sequences

### v1

@@ Appendix B Evaluation Benchmarks
Appendix C Experiment Details
C.1 Baseline Comparison
Table 6 presents the time per sample for vanilla autoregressive generation, Ring Attention, and Star Attention across sequence lengths ranging from 16K to 128K. The results indicate that both Ring and Star Attention can process sequences up to 128K tokens on 8 A100 GPUs, whereas vanilla autoregressive inference encounters OOM issues beyond 64K tokens. For sequence lengths below 32K, vanilla inference is faster than the distributed attention mechanisms, primarily due to the GPU communication overhead incurred in the distributed setups. However, in long context scenarios i.e. on sequence lengths exceeding 32K tokens, Star Attention begins to demonstrate clear performance advantages. As demonstrated in Table 1, the speedup achieved by Star Attention increases significantly with longer sequence lengths.
Seq. Length Time Per Sample (s)
(K) Vanilla Ring Star
16 7 10 9
32 10 12 10
64 18 22 12
128 OOM 53 20
Table 6: Time per sample (seconds) for Llama3.1-8B-Instruct model with vanilla (global) inference, ring (global) and star attention, using 8 A100 GPUs. Vanilla autoregressive generation encounters out-of-memory (OOM) at 128K sequence length. It performs best in short context scenarios (i.e. sequences upto 32K tokens) but in long context scenarios, star attention demonstrates significant speedup.
C.2 Hardware for Inference Speed
Model Size Seq. Length # GPUs # Workers
8B 16K - 128K 8 4
256K - 512K 16 8
1M 32 16
70B 16K - 32K 8 4
64K 16 4
128K 32 8
Table 7: Resources used for the speedup experiments
C.3 Prompt Templates
Prompt structure for base models:
1{context}{query}{answer_prefix}
Prompt structure for Llama-3 and Llama-3.1 Instruct models:
2
4
6
7{answer_prefix}
Appendix D Accuracy of Star Attention
Table 8 shows the exact accuracy scores of star attention vs global attention across the RULER and the BABILong benchmark from Figure 4.
Model Seq. Block RULER (%) BABILONG (%)
length size Global Star $\Delta$ Global Star $\Delta$
GradientAI Llama-3-8B -Instruct-262k 16K 4K 88.92 89.48 +0.63% 43.60 43.40 -0.46%
32K 8K 85.25 85.74 +0.58% 40.00 39.40 -1.50%
64K 16K 83.17 82.30 -1.05% 40.40 39.00 -3.47%
128K 32K 79.25 77.79 -1.83% 30.80 33.20 +7.79%
Meta Llama-3.1-8B -Instruct 16K 4K 99.78 91.27 -1.02% 59.60 59.80 +0.34%
32K 8K 99.66 88.70 +1.34% 54.60 54.00 -1.10%
64K 16K 98.72 83.37 -1.67% 49.20 46.60 -5.28%
128K 32K 92.54 74.41 -2.49% 40.00 38.60 -3.50%
Meta Llama-3.1-8B -Base 16K 4K 77.18 78.64 +1.9% 22.00 25.20 +14.55%
32K 8K 74.76 76.91 +2.88% 22.60 24.00 +6.19%
64K 16K 70.01 69.09 -1.32% 26.80 27.20 +1.49%
128K 32K 64.68 69.58 +7.58% 31.00 26.40 -14.84%
Table 8: Accuracy (%) of star attention on RULER and BABILONG evaluated on sequence lengths of 16K, 32K, 64K, and 128K. In all experiments, the block size and anchor block size are set to one-quarter of the total sequence length. Results using the Llama-3-8B-Instruct-262k, Llama-3.1-8B-Instruct and Llama-3.1-8B-Base models demonstrate that star attention retains 95-100% of the accuracy of global attention, and in some cases, even outperform it.
Appendix E Accuracy on all RULER Tasks
This section contains the accuracy of all the models we evaluated across all 13 tasks in RULER.
Llama-3.1-8B-Instruct
Block Size (K) Seq. Len. (K) Retrieval (NIAH)
Single 1 Single 2 Single 3
Multi-
Key 1
Multi-
Key 2
Multi-
Key 3
Multi-
Value
Multi-
Query
Global Attn. 16 100 100 100 99.8 100 99 99.9 99.5
32 100 100 100 99.8 99.8 99.6 99 99.05
64 100 100 100 99.4 99.2 96.8 95.15 99.2
128 100 99.6 99.8 97.6 87.2 66.8 91.55 97.8
4 16 100 99.4 100 98 98.8 99 91.1 98.25
8 32 100 100 100 99.2 99.4 98.2 94 98.3
16 64 100 100 100 99.2 98 90 85.35 97.9
32 128 100 100 99.6 96.4 84.8 59 82.7 96.55
Table 9: Accuracy of Llama-3.1-8B-Instruct on retrieval tasks in RULER with Global Attention and Star Attention
Llama-3.1-8B-Instruct
Multi-Hop Aggregation Question Answering
Block Size (K) Seq. Len. (K) VT CWE FWE QA 1 QA 2
Global Attn. 16 99.56 75 88.87 80.8 56.4
32 99.2 14.7 93.93 78.8 54
64 95.44 1.96 85.13 78.8 51.2
128 61.76 0.04 72.33 76 41.6
4 16 91.96 85.72 89.73 80.2 54.4
8 32 92.68 45.66 95.27 78.6 51.8
16 64 92.32 5.78 86.47 78.4 50.4
32 128 62.8 0.04 75.87 68 41.6
Table 10: Accuracy of Llama-3.1-8B-Instruct on multi-hop, aggregation, and question answering tasks in RULER with Global Attention and Star Attention
Llama-3.1-8B-Base
Block Size (K) Seq. Len. (K) Retrieval (NIAH)
Single 1 Single 2 Single 3
Multi-
Key 1
Multi-
Key 2
Multi-
Key 3
Multi-
Value
Multi-
Query
Global Attn. 16 100 100 100 99.2 100 99.4 99.45 99.85
32 100 100 100 99 99.4 99.4 99.55 99.4
64 100 100 100 98.8 86.2 95.4 96.8 97.55
128 100 100 98 93.8 53.6 64 80.9 85.3
4 16 100 100 100 97.4 99.2 99 98.4 99.15
8 32 100 100 100 96.2 98.2 99.2 98.55 98.7
16 64 100 100 100 96.6 90.6 85.6 94.9 96.15
32 128 100 100 98.2 88.8 67 47.6 72.75 77.55
Table 11: Accuracy of Llama-3.1-8B-Base on retrieval tasks in RULER with Global Attention and Star Attention
Llama-3.1-8B-Base
Multi-Hop Aggregation Question Answering
Block Size (K) Seq. Len. (K) VT CWE FWE QA 1 QA 2
Global Attn. 16 99.92 65.66 17.4 11 11.4
32 99.28 23.56 28.27 13.8 10.2
64 96.8 2.04 13.73 14.2 8.6
128 71.68 0.64 30.53 51.2 11.2
4 16 97.24 86.46 20.67 11.6 13.2
8 32 97.2 58.72 30.47 11.8 10.8
16 64 94.44 8.86 11.2 10.6 9.2
32 128 81.6 2.98 81.27 48.2 38.6
Table 12: Accuracy of Llama-3.1-8B-Base on multi-hop, aggregation, and question answering tasks in RULER with Global Attention and Star Attention
Llama-3.1-70B-Instruct
Block Size (K) Seq. Len. (K) Retrieval (NIAH)
Single 1 Single 2 Single 3
Multi-
Key 1
Multi-
Key 2
Multi-
Key 3
Multi-
Value
Multi-
Query
Global Attn. 16 100 100 100 97.8 99.8 98.6 99 99.65
32 100 100 100 99.6 99 99 99.1 100
64 100 100 100 99.8 96 97.6 95.65 99.95
128 97.2 99.2 99.4 93 26 27.4 92.05 92.45
4 16 100 100 100 97.4 99.4 99.2 80.9 97.2
8 32 100 100 100 98.2 96.4 95 87.85 97.4
16 64 100 100 100 98 93.2 95.4 86.25 96.45
32 128 98.2 100 99.4 80 19.2 16.4 61.65 70.8
Table 13: Accuracy of Llama-3.1-70B-Instruct on retrieval tasks in RULER with Global Attention and Star Attention
Llama-3.1-70B-Instruct
Multi-Hop Aggregation Question Answering
Block Size (K) Seq. Len. (K) VT CWE FWE QA 1 QA 2
Global Attn. 16 100 99.3 97 82.6 62.4
32 100 94.22 98.87 80.4 59.8
64 100 39.7 93.73 74.6 54
128 50.08 2.98 77 58.4 33.6
4 16 87.32 99.52 97.13 82.2 60.6
8 32 90.08 94.5 99.2 80.2 58
16 64 91.52 49.54 94.93 73.4 53.6
32 128 41.4 2.7 80.07 50.8 31
Table 14: Accuracy of Llama-3.1-70B-Instruct on multi-hop, aggregation, and question answering tasks in RULER with Global Attention and Star Attention
GradientAI Llama-3-8B-Instruct-262K
Block Size (K) Seq. Len. (K) Retrieval (NIAH)
Single 1 Single 2 Single 3
Multi-
Key 1
Multi-
Key 2
Multi-
Key 3
Multi-
Value
Multi-
Query
Global Attn. 16 100 100 99.8 99.6 100 96 95.35 99.85
32 100 100 100 99.8 100 95 96.2 99.75
64 100 100 100 98.4 99.4 91.4 97.75 99.6
128 100 97.8 98.8 98.8 99.8 79.8 94.65 98.05
256 100 100 99.4 96.4 89.6 25.6 87.3 93.2
4 16 100 98.4 96.6 99.6 99.4 97 89.2 99.75
8 32 100 100 100 99.2 99.6 96 91.6 99.7
16 64 100 100 100 99.4 99.4 90 91.45 99.3
32 128 100 100 100 98.4 97.8 66.8 89.3 96.8
32 256 100 99.6 98.4 91.4 53 23 75 81.05
Table 15: Accuracy of GradientAI Llama-3-8B-Instruct-262K on retrieval tasks in RULER with Global Attention and Star Attention
GradientAI Llama-3-8B-Instruct-262K
Multi-Hop Aggregation Question Answering
Block Size (K) Seq. Len. (K) VT CWE FWE QA 1 QA 2
Global Attn. 16 95.36 42.1 91.07 80.2 56.6
32 93.88 4.5 90.53 74 54.6
64 92.28 0.22 82.73 69.8 49.6
128 77.88 0.36 73.27 65.6 45.4
256 52.8 1.8 77.93 67 37
4 16 90.64 67.32 90.53 77.2 57.6
8 32 92.16 19.2 89.6 73.8 53.8
16 64 88.6 0.4 84.13 69.6 47.6
32 128 81.12 0.3 75.4 61.6 43.8
32 256 72.64 1.8 81.6 61.6 33
Table 16: Accuracy of GradientAI Llama-3-8B-Instruct-262K on multi-hop, aggregation, and question answering tasks in RULER with Global Attention and Star Attention
GradientAI Llama-3-8B-Instruct-1048K
Block Size (K) Seq. Len. (K) Retrieval (NIAH)
Single 1 Single 2 Single 3
Multi-
Key 1
Multi-
Key 2
Multi-
Key 3
Multi-
Value
Multi-
Query
Global Attn. 16 100 99.2 100 99 99.6 90.2 96.1 99.25
32 100 100 100 99.4 99.2 69.8 96.3 98.45
64 100 100 100 99 99 51.4 96 98.75
128 100 98.2 99.8 99.8 98.8 42.8 98.2 97.75
256 100 100 100 98.4 97 22.4 96.1 97.15
512 100 99.8 100 95.6 88.4 9.4 89.25 92.55
1024 99.4 99.4 100 92.6 67.8 1.4 82 88.85
4 16 100 98 96.8 98.6 99 94.4 90.3 98.1
8 32 100 99.8 100 98.8 99.4 87 91 97.6
16 64 100 100 100 99.4 99 66.8 92.3 97.95
32 128 100 100 100 99.4 98.4 62.8 92.25 96.8
32 256 100 99.8 100 95.4 90.4 53.8 76.5 88.6
32 512 99.8 95.8 97.6 85.8 64.2 19.4 57.2 63.8
32 1024 99.6 97.2 100 84.2 27 1 55.15 60.3
Table 17: Accuracy of GradientAI Llama-3-8B-Instruct-1048K on retrieval tasks in RULER with Global Attention and Star Attention
GradientAI Llama-3-8B-Instruct-1048K
Multi-Hop Aggregation Question Answering
Block Size (K) Seq. Len. (K) VT CWE FWE QA 1 QA 2
Global Attn. 16 93.32 22.52 88.73 76.8 54.8
32 91.96 0.54 87.27 75.4 54.4
64 81.6 0.32 79.8 75.4 46.4
128 76.68 0.22 76.67 68.4 48.8
256 63 0.22 78.27 71.4 43.8
512 34.8 0.86 85.6 66.8 37.8
1024 24.28 3.5 72.53 66.2 30.2
4 16 90 62.64 86.53 77.4 55.4
8 32 90.08 11.38 85.2 75.2 53.8
16 64 81.96 0.38 82 74.4 46.6
32 128 79.32 0.22 78.33 64.6 46.4
32 256 66.4 0.22 81.53 67.4 37.6
32 512 50.76 0.24 85 63 30.8
32 1024 63.68 6.9 79.67 57.8 26.4
Table 18: Accuracy of GradientAI Llama-3-8B-Instruct-1048K on multi-hop, aggregation, and question answering tasks in RULER with Global Attention and Star Attention
Figure 8: Accuracy (%) of star attention using the Llama-3.1-8B-Instruct model on the 5 categories of tasks in RULER on sequence lengths of 16K, 32K, 64K, and 128K. In all experiments, the block size and anchor block size are set to one-quarter of the total sequence length. For the NIAH and QA tasks, Star Attention retains upto 95-100% accuracy of the baseline. The Multi-Hop Tracing task is notably challenging because it requires inter-block communication, which leads to expected performance degradation. Interestingly, Star Attention performs better with sequence lengths of 128k on this task, but this may be due to noise given the suboptimal baseline. In aggregation tasks, Star Attention show significant improvement as distributed local attention helps the model in such summarization tasks.

### v2

@@ Appendix C Evaluation Benchmarks
InfiniteBench: This benchmark comprises 10 real-world and synthetic tasks, each crafted to assess different aspects of language processing and comprehension in extended contexts. Details of each task is shown in 10
Avg. input Avg. output
Task name Context # samples tokens tokens
En.Sum Fake Book 103 171.5k 1.1k
En.QA Fake Book 351 192.6k 4.8
En.MC Fake Book 229 184.4k 5.3
En.Dia Script 200 103.6k 3.4
Zh.QA New Book 175 2068.6k 6.3
Code.Debug Code Document 394 114.7k 4.8
Code.Run Synthetic 400 75.2k 1.3
Math.Calc Synthetic 50 43.9k 43.9k
Math.Find Synthetic 350 87.9k 1.3
Retrieve.PassKey Synthetic 590 122.4k 2.0
Retrieve.Number Synthetic 590 122.4k 4.0
Retrieve.KV Synthetic 500 89.9k 22.7
Table 10: Configuration of tasks in InfiniteBench
Appendix D RULER Analysis
Figure 7: Accuracy (%) of star attention using the Llama-3.1-8B-Instruct model on the 5 categories of tasks in RULER on sequence lengths of 16K, 32K, 64K, and 128K. In all experiments, the block size and anchor block size are set to one-quarter of the total sequence length. For the NIAH and QA tasks, Star Attention retains upto 97-100% accuracy of the baseline. The Multi-Hop Tracing task is notably challenging because it requires inter-block communication, which leads to expected performance degradation. Interestingly, Star Attention performs better with sequence lengths of 128k on this task, but this may be due to noise given the suboptimal baseline. In aggregation tasks, Star Attention show significant improvement as distributed local attention helps the model in such summarization tasks.